lowing is a set of data for a small company. Using this data, analyze the expenses. Using the functions in the spreadsheet, calculate the monthly average for each expense for the year Using the calculations, calculate the percentage of each expenses to the total sales Feb- Mar- May- 14 Aug- Sep- 14 14 Jan-14 14 Jun-14 Jul-14 14 Dec-1 Sales Revenue Actual Actual Actual Actual Actualal Actualcal ctal Actual Actual Actual Product 1 48267 50,680 52708 55,343 57,003 56,433 56456 57,020 57,305 57,549 57,995 58,343 Product 2 10,214 10,725 | 11,15411,711 12,063 11,942 1242 12,792 13,432 13,513 13,594 13,675 Product 3 5,845 6137 6,383 6,386 6388 6324 6577 6774 6778 6782 6,788 6,829 Total Sales 64,326 | 67,542 | 70,244 | 73,440 | zsAsa | 74,699丨25-25 sez | ZZS15 | ZZ944 nventory 20,584 20,263 24 585 23,501 24,145 22410 24245 28,337 21,704 24942 25081 25.23 Supplies 3,216 4053 3,512 3,672 6,036 3.735 5,282 2.298 3,101 3,89 3919 3.942 Direct 12,865 13,508 10,537 14.688 13.582 18675 15,090 19,913 15,503 16.368 15,675 15.770 Total Cost of Goods 36,666 37,824 38,634 41,861 43,763 44,820 44517 50,547 40,308 45,207 4,675 44943 Salaries 12.879 13,179 12.879 1279 1288 12780 1245 12250 12.950 12963 12.850 12 100 Utilities 780 | 841 | 740 | 685 | 710 | 795 | 910 180 | 1,020|650 710 745 Operating Expenses Advertising 400 410 1,150 1,250 1,200 1,000 Insurance 1,522 1,522 1,522 1,522 1,522 1,522 1,522 1,522 1,522 1,522 1,522 1,522 100 741 800 280 160 40 Total Operating Expenses 15,741 15,542 15,141 15,326 15,114 15,838 16,077 15,362 16,642 16,385 16,562 15,467 EBIT-一ㄧ 27,660 29,719 31,610 31,579|31 6ll 29,880 3a936-040 322n2-236 33702-0- 9681 10,402 11,063 11,053 11,092 10,458 10827 9,114 13,023 11,458 11,796 11,867 1,252 | 1,252 | 1,2sa l 1,252 1,252 ↓ 1,252 1,2 1,252 1,252 ncome 16,727 18,065 19,294 19,275 19.347 18,170 18856 15.674 22 933 20,022 20.654 20 286 ncome Taxes Interest 252
Answer #1

Monthly Percentage Average to total sales Jan 14 Feb 14 Mar 14 Apr 14 May 14 Jun 14 Jul 14 Aug 14 Sep 14 Oct 14 Nov 14 Dec 14 2 Sales Revenue 3 Product 1 4 Product 2 5 Product 3 6 Total Sales 7 Direct Expenses 8 Inventory 9 Labor 10 Supplies 11 Total cost of goods 36,665 37,824 38,634 41,861 43,763 44,820 44,517 50,548 40,308 45,207 44,675 44,943 42,814 12 Operating Expenses 13 Salaries 14 Advertising 15 Utilities 16 Insurance 17 Repairs 48,26750,680 52,708 55,343 57,003 56,433 56,456 57,020 57,305 57,649 57,995 58,343 10,214 10,725 11,154 11,711 12,063 11,942 12,420 12,792 13,432 13,513 13,594 13,675 5,845 6,137 6,383 6,386 6,388 6,3246,577 6,774 6,778 6,782 6,788 6,829 64,326 67,54270,245 73,440 75,454 74,699 75,45376,586 77,515 77,944 78,377 78,847 890,428 32% 20,584 20,26324,585 23,501 24,145 22,410 24,145 28,337 21,704 24,942 25,081 25,231 23,744 12,865 13,508 10,537 14,688 13,582 18,675 15,090 19,913 15,503 16,368 15,675 15,770 15,181 3,216 4,053 3,512 3,672 6,036 3,735 5,282 2,298 3,101 3,897 3,919 3,942 | 3,889 5% 58% 12,879 13,179 12,879 12,579 12,882 12,780 12,845 12,250 12,950 12,963 12,850 12,100 12,761 410 1,150 1,250 1,200 1,000 493 780 841 740 685 710 795 910 1,180 1,020 650 710 745 814 1,522 1,522 1,522 1,522 1,522 1,522 1,522 522 1,522 ,522 1,522 1,5221,522 280 100177 17% 196 196 2% 400 500 160 40 741 800 Total Operating 18 Expenses 19 EBIT 20 Income Taxes 21 Interest Expense 1,252 1,252 1,252 1,252 1,252 1,252 1,252 1,252 1,252 1,252 1,252 1,2521,252 22 Net Income | 21% 15,741 15,542 15,141 15,326 15,114 15,838 16,077 15,362 16,642 16,38516,562 15,467 15,766 11,920 14,176 16,470 16,253 16,577 14,041 14,859 10,676 20,565 16,352 17,140 18,437 9,681 10,402 11,063 11,053 11,092 10,458 10,827 9,114 13,023 11,458 11,796 11,867 10,986 15% 2% 987 2,522 4,155 3,948 4,233 2,331 2,780 310 6,290 3,642 4,092 5,318

formula:

0 Feb 14 Mar 14 Apr 14 May 14 Aug 14 Oct 14 Nov 14 Dec 14 Monthly AverogePercentoge to total sales 3 4 5 6 Product 1 48267 Product 210214 Product 3 5845 Total Sales -SUM(B3B5) 57649 11942 12792 13594 13675 =SUM(C3CS)-SUMID3DS rSUM(E3E5)-SUN1(F3FS)-SUM(G3 G5)-SUM(H3HS)-SUM(13:15) =SUM(J35) :SUM(K3K5)-SUM(し3:L5) -SUM(M3 MS) SUM(B6:M6) 23501 24145 22410 21704 8 Inventory 20584 9 Labor 10 Supplies 3216 24942 25081 3672 SUMB10M0)/12-SUMB10 MIO/SOS6 Total cost of «SUMIB8 В 10 -SUM(C8:C10-SUMID8 DI-SUM(Ε8:E10-SUM(F8:F 10-SUM(G8 G 10-SUM(H8:H 10-SUM(18 110)-SUM(18:Ј10)-SUM(K8:K10-SUM(L8:L10 -SUM(M8:M10) -SUMB11:M11)/12 -SUM B11:M11) SOS6 13 Slaies 12879 4 Advertising 400 15 Utilities 16 Insurance 1522 17 Repairs 12780 -SUMB13 M13)/12-SUMB13 M13/SOS6 SUMB14 M14/12SUMB14 M14/SOS6 SUMB15:15/12-SUMB15 M15)/SOS6 -SUM B16 M16)/12-SUM(B16 MI6/SOS6 -SUMIB17M1 7)/12 -SUMB17 M17/SOs6 Total Operating Expenses 18 SUM(B13 B1 -SUM(C13 CI-SUMID13 D -SUMIE 13 El-SUM(F13:F1-SUM(G13 Gİ-SUM(H13 HI-SUM(113 11:-SUM J13:11:-SUMIK13:K1-SUM L13:L1 -SUM(M13 M17-SUM B18 M18 /12 -SUMB18 M18) SOS6 :86-811-818 弍6-C11-C18 ะD6-D11-D18=E6-E11-E18 zf6-F11-F18 =G6-G11-G18=H6-Н 11-H18=16-111-118 =J6-J11-J18 =KS-K11-K18 =L6-し11-L18 =M6-М 11-M 18 11092 10827 11458 11796 SUM B20 M20)/12-SUMB20:M20)/SOS6 21 22 Expense Net Income 1252 -819-B20-B21弍19-C20-C21=D19-D20-D2=E19-E20-E21=F19-F20-F21=G19-G20-G21=H 19-H20-H21=119-120-121 SUM B21 M21)/12-SUMB21 M21) SOSe =」19-120-121 -K19-K20-K21-L19-L20-L21:M19-M20-M21